# Variability study of classical supergiant X-ray binary 4U 1907+09 using NuSTAR

### v2

@@ 4 Discussion
4U 1907+09 is a classical supergiant X-ray binary known to show the dips and flares in its light curve. In this work, we analyzed the NuSTAR data of 4U 1907+09 observed during November 2024. The spin down/spin up and torque reversals of the pulsar were documented in several studies (Cook and Page, 1987; in ’t Zand et al., 1998; Baykal et al., 2001, 2006; Mukerjee et al., 2001; Fritz et al., 2006; Inam et al., 2009; Şahiner et al., 2012; Varun et al., 2019; Tobrej et al., 2023). In the most recent study, Tobrej et al. (2023) found that the 4U 1907+09 has been spinning down. In this work, we estimated the pulse period of 4U 1907+09 to be $\sim 443.99(4)$ s. This indicates that the source is still spinning down. The pulse profile is one of the indicators of the emission region from the neutron star. A double-peaked (at $\sim 0.26$ and $0.73$) asymmetric pulse shape was observed in 4U 1907+09, as shown in the left panel of Figure 3. A similar pulse shape was also reported for the 4U 1907+09 in the previous studies (Maitra and Paul, 2013; Varun et al., 2019; Tobrej et al., 2023). The observed asymmetric pulse shape is most likely caused by a magnetic dipole offset from the rotational axis of the pulsar (Parmar et al., 1989; Leahy, 1991; Riffert et al., 1993; Bulik et al., 1995). Another plausible explanation for the asymmetric shape of pulse profiles is an asymmetric accretion stream (Basko and Sunyaev, 1976; Wang and Welter, 1981; Miller, 1996). As energy increases, the pulse shape changes significantly. The peak at 0.26 begins to fade at 15 keV. The PF followed an increasing behaviour with energy.
@@
The NuSTAR spectrum was fitted with three different models: power law with highecut, compTT, and compmag. Table 1 contains the corresponding fit parameters. The iron line and CRSF with its harmonic were found at $\sim 6.31$ keV, $\sim 17.6$ keV, and $\sim 38.02$ keV, respectively. After incorporating gravitational redshift, the CRSF corresponds to a magnetic field of $\sim 2\times 10^{12}$ G. The average flux of the source during the on-state in the 3–50 keV energy range was found to be $\sim 6.59\times 10^{-10}\mathrm{erg\;cm^{-2}\;s^{-1}}$. Assuming a distance of 1.9 kpc, the estimated luminosity of the source is $\sim 2.85\times 10^{35}\mathrm{erg\;s^{-1}}$. The luminosity of the source is similar to that observed by NuSTAR in 2018 (Tobrej et al., 2023). Using the formalism in Becker et al. (2012), we estimate the critical luminosity $L_{\mathrm{crit}}=1.5\times 10^{37}\;\left(\frac{B}{10^{12}\;\mathrm{G}}\right)^{16/15}\;\mathrm{erg\;s^{-1}}$ to be $3\times 10^{37}\mathrm{erg\;s^{-1}}$. The Coulomb braking luminosity $L_{\mathrm{coul}}=1.2\times 10^{37}\;\left(\frac{B}{10^{12}\;\mathrm{G}}\right)^{1/3}\;\mathrm{erg\;s^{-1}}$ to be $1.5\times 10^{37}\mathrm{erg\;s^{-1}}$ (Becker et al., 2012; Langer and Rappaport, 1982). The observed luminosity of 4U 1907+09 is much lower than the $L_{\mathrm{crit}}$ and $L_{\mathrm{coul}}$. As a result, radiation from a gas-mediated shock without collision is anticipated. This leads to a ”pencil” beam, where the radiation is emitted mostly along the magnetic field lines.
@@
Numerous studies examine the dips and flares in the light curve of 4U 1907+09 (in ’t Zand et al., 1997; Doroshenko et al., 2012; Şahiner et al., 2012; Shakura et al., 2012; Varun et al., 2019). Apart from 4U 1907+09, off-states or dips have been reported in the Vela X-1 (Kreykenbohm et al., 2008; Doroshenko et al., 2011) and GX 301-2 (Göğüş et al., 2011). Previous studies have ruled out the likelihood of dips generated by a dense, clumpy wind obscuring the neutron star (Fürst et al., 2011; Şahiner et al., 2012). Shakura et al. (2012) suggested that the onset of the off-state in these systems is due to the transition from the Compton cooling-dominated to the radiative cooling-dominated regime. During this observation, the source showed the dip and flare in the light curve. During the dip, the flux of the source decreased 8 times the flux during the low flux on-state. The flux during high flux on-state increased 3 times during low flux on-state. As the flux varied during the on-state, the pulse profile did not differ much and followed a similar trend (as shown in the right panel of Figure 3). It indicates that the emission process did not change during the flare activity. The spectral modelling of 4U 1907+09 revealed that the $E_{\mathrm{CRSF}}$ is unaltered at $\sim 18$ keV in both the high flux (flare) on-state and the off-state. It suggests that the accretion column may not be formed. The e-folding energy during the high flux on-state is lower than the low flux high state. It shows that the high flux on-state is softer than the low flux on-state. The line width of the CRSF during the high flux on-state is higher than the low flux on-state.
@@ 5 Summary
    The CRSF at $\sim 18$ keV was found in both on and off states. The width of CRSF correlated positively with the flux.
@@ Acknowledgments
We have utilized the archived NuSTAR data provided by the High Energy Astrophysics Science Archive Research Center (HEASARC) online service maintained by the Goddard Space Flight Center. For this work, we have made use of the NuSTAR Data Analysis Software (NuSTARDAS) jointly developed by the ASI Space Science Data Center (SSDC, Italy) and the California Institute of Technology (Caltech, USA). RK acknowledges the Department of Astronomy and Astrophysics, Tata Institute of Fundamental Research, Mumbai, India for providing research facilities.

### v3

@@ 4 Discussion
4U 1907+09 is a classical sgXB known to show dips and flares in its light curve. In this work, we analyzed the NuSTAR data of 4U 1907+09 observed during November 2024. The spin down/spin up and torque reversals of the pulsar were documented in several studies (Cook and Page, 1987; in ’t Zand et al., 1998; Baykal et al., 2001, 2006; Mukerjee et al., 2001; Fritz et al., 2006; Inam et al., 2009; Şahiner et al., 2012; Varun et al., 2019; Tobrej et al., 2023). In the most recent study, Tobrej et al. (2023) found that 4U 1907+09 has been spinning down. In this work, we estimated the pulse period of 4U 1907+09 to be $\sim 443.99(4)$ s. This indicates that the source is still spinning down. At low or moderate luminosities, a hot quasi-static shell develops as captured wind matter settles subsonically on the NS magnetosphere (Davies and Pringle, 1981). During spin-down episodes, this shell removes angular momentum from the rotating NS magnetosphere through large-scale convective processes Shakura et al. (2012) and could be responsible for 4U 1907+09’s long-term spin-down behaviour.
@@
The pulse profile is one of the indicators of the emission region from the neutron star. A double-peaked (at $\sim 0.26$ and $0.73$) asymmetric pulse shape was observed in 4U 1907+09, as shown in the left panel of Figure 3. A similar pulse shape was also reported for the 4U 1907+09 in the previous studies (Maitra and Paul, 2013; Varun et al., 2019; Tobrej et al., 2023). The observed asymmetric pulse shape is most likely caused by a magnetic dipole offset from the rotational axis of the pulsar (Parmar et al., 1989; Leahy, 1991; Riffert et al., 1993; Bulik et al., 1995). Another plausible explanation for the asymmetric shape of pulse profiles is an asymmetric accretion stream (Basko and Sunyaev, 1976; Wang and Welter, 1981; Miller, 1996). As energy increases, the pulse shape changes significantly. The peak at 0.26 begins to fade at 15 keV. For Vela X-1, Shakura et al. (2013) revealed that when the source flips from on-state to off-state, it transitions from a fan beam pattern to a pencil beam pattern in the 20–60 keV energy range. Furthermore, Vela X-1 exhibited absorption in the pencil beam at lower energies. However, 4U 1907+09 displayed the pencil beam during the on-state, without indicating a fan beam pattern. Furthermore, no absorption was detected in the pencil beam at low energies. It indicates that the optical depth of accretion flow above the polar cap is always less than 1 in 4U 1907+09 during the transition from on-state to off-state. The PF of X-ray pulsars shows the complex behaviour and dips at cyclotron energy (Lutovinov and Tsygankov, 2009; Ferrigno et al., 2023). Tobrej et al. (2023) observed a dip near the cyclotron energy in PF. The PF in our investigation showed increasing behaviour with energy and did not exhibit a complicated structure because of the limited number of data points.
@@
The NuSTAR spectrum was fitted with three different models: power law with highecut, compTT, and compmag. Table 1 contains the corresponding fit parameters. The iron line and CRSF with its harmonics were found at $\sim 6.31$ keV, $\sim 17.6$ keV, and $\sim 38.02$ keV, respectively. After incorporating gravitational redshift, the CRSF corresponds to a magnetic field of $\sim 2\times 10^{12}$ G. The average flux of the source during the on-state in the 3–50 keV energy range was found to be $\sim 6.59\times 10^{-10}\mathrm{erg\;cm^{-2}\;s^{-1}}$. Assuming a distance of 1.9 kpc, the estimated luminosity of the source is $\sim 2.85\times 10^{35}\mathrm{erg\;s^{-1}}$. The luminosity of the source is similar to that observed by NuSTAR in 2018 (Tobrej et al., 2023). Using the formalism in Becker et al. (2012), we estimate the critical luminosity $L_{\mathrm{crit}}=1.5\times 10^{37}\;\left(\frac{B}{10^{12}\;\mathrm{G}}\right)^{16/15}\;\mathrm{erg\;s^{-1}}$ to be $3\times 10^{37}\mathrm{erg\;s^{-1}}$. The Coulomb braking luminosity $L_{\mathrm{coul}}=1.2\times 10^{37}\;\left(\frac{B}{10^{12}\;\mathrm{G}}\right)^{1/3}\;\mathrm{erg\;s^{-1}}$ to be $1.5\times 10^{37}\mathrm{erg\;s^{-1}}$ (Becker et al., 2012; Langer and Rappaport, 1982). The geometry of the accretion column and emission beam pattern changes with the luminosity. At high luminosity ($L_{X}\sim 10^{37-38}\;\mathrm{erg\;s^{-1}}$), the radiation predominantly escapes through the column walls in the sinking zone, resulting in a ”fan beam” (Davidson, 1973). At low luminosity ($L_{X}\sim 10^{35}\;\mathrm{erg\;s^{-1}}$), the infalling matter impacts the NS surface, and radiation is released from the top of the column, resulting in a ”pencil beam” pattern (Burnard et al., 1991; Nelson et al., 1993). In the intermediate range, $L_{X}\sim 10^{35-37}\;\mathrm{erg\;s^{-1}}$, the radiation emits from the wall as well as the top of the accretion column, resulting in a complicated emission pattern composed of fan beams and pencil beams (Blum and Kraus, 2000). The observed luminosity of 4U 1907+09 is much lower than the $L_{\mathrm{crit}}$ and $L_{\mathrm{coul}}$. As a result, radiation from a gas-mediated shock without collision is anticipated. This leads to a ”pencil” beam, where the radiation is emitted mostly along the magnetic field lines.
@@
Numerous studies examine the dips and flares in the light curve of 4U 1907+09 (in ’t Zand et al., 1997; Doroshenko et al., 2012; Şahiner et al., 2012; Shakura et al., 2012; Varun et al., 2019). Apart from 4U 1907+09, off-states or dips have been reported in the Vela X-1 (Kreykenbohm et al., 2008; Doroshenko et al., 2011) and GX 301-2 (Göğüş et al., 2011). Previous studies have ruled out the likelihood of dips generated by a dense, clumpy wind obscuring the neutron star (Fürst et al., 2011; Şahiner et al., 2012). Shakura et al. (2012) suggested that the onset of the off-state in these systems is due to the transition from the Compton cooling-dominated to the radiative cooling-dominated regime. During this observation, the source showed a dip and flare in the light curve. During the dip, the flux of the source decreased 8 times the flux during the low-flux on-state. The flux during high flux on-state increased 3 times during low flux on-state. As the flux varied during the on-state, the pulse profile did not differ much and followed a similar trend (as shown in the right panel of Figure 3). This indicates that the emission process did not change during the flare activity. The cyclotron energy correlates with the luminosity (Staubert et al., 2019). In the super-critical state, the cyclotron energy displays a negative correlation with the luminosity. The cyclotron energy remains constant or has a positive correlation with luminosity in the sub-critical phase. The spectral modelling of 4U 1907+09 revealed that the $E_{\mathrm{CRSF}}$ is unaltered at $\sim 18$ keV in both the high flux (flare) on-state and the off-state. It suggests that the accretion column may not be formed. The off-state spectrum is softer than the on-state spectrum. This behaviour is also seen in GX 301-2 (Göğüş et al., 2011) and Vela X-1 (Doroshenko et al., 2011). The e-folding energy during the high flux on-state is lower than the low flux high state. The line width of the CRSF during the high flux on-state is higher than the low flux on-state.
In addition, the polarimetric studies of strongly magnetized X-ray pulsars are useful for the determination of XRP geometry and related astrophysical phenomena (Meszaros et al., 1988; Santangelo et al., 2019; Caiazzo and Heyl, 2021; Poutanen et al., 2024). 4U 1907+09 is an ideal candidate to study the polarimetric properties in the subcritical accretion regime. For 4U 1907+09, the vacuum polarization effects are anticipated to be important (Shakura et al., 2013). For Vela X-1 (a wind-fed sgXB similar to 4U 1907+09), the Imaging X-ray Polarimetry Explorer (IXPE; Weisskopf et al. (2022)) found a polarization degree (PD) of 2.3% ± 0.4% at the polarization angle (PA) of -47.3 ± 5.4 (Forsblom et al., 2023). Finally, we can speculate a small PD for 4U 1907+09 as detected in Vela X-1.
@@ 5 Summary
    The CRSF at $\sim 18$ keV was found in both on and off states. While the CRSF energy remained constant, CRSF’s width showed a positive correlation with flux.
@@ Acknowledgments
We thank the anonymous referee for their comments and recommendations, which enhanced the quality of this paper. We have utilized the archived NuSTAR data provided by the High Energy Astrophysics Science Archive Research Center (HEASARC) online service maintained by the Goddard Space Flight Center. For this work, we have made use of the NuSTAR Data Analysis Software (NuSTARDAS) jointly developed by the ASI Space Science Data Center (SSDC, Italy) and the California Institute of Technology (Caltech, USA). RK acknowledges the Department of Astronomy and Astrophysics, Tata Institute of Fundamental Research, Mumbai, India, for providing research facilities.